OpenStudy (anonymous):

Solve the equations and check your solutions. If there is no solution ,say so 1/x+2/3x=1/3 need the steps please.

OpenStudy (anonymous):

\[\frac{1}{x} + \frac{2}{3x} = \frac{1}{3}\] \[\frac{3}{3x} + \frac{2}{3x} = \frac{1}{3}\] \[\frac{5}{3x} = \frac{1}{3}\] \[3x=15\] \[x=5\]

OpenStudy (anonymous):

Thanks a lot